OpenStudy (anonymous):

Solve the quadratic equation x^2 − 4x + 57 = -5 by completing the square. *medals given*

OpenStudy (anonymous):

i took (x^2 -4x) = -62 then complete square x^2 - 4x +4 = -62 +4

OpenStudy (anonymous):

i got (x-2)^2 = -58 then x-2 = plus minus sqrt-58

OpenStudy (anonymous):

on wolfram site says answer would be 2 + i sqrt58 , 2 - i sqrt58